OpenStudy (anonymous):

.Which of the following equations has no real solutions? A. 9x - 3 = 9x - 3 B. 9(x - 3) = 9x - 27 C. 9x - 3 = 27x - 3 D. 9(x + 3) = 9(x + 11)

OpenStudy (anonymous):

@armyengineer225 can u help me on this one?

OpenStudy (anonymous):

First simplify each equations\

OpenStudy (anonymous):

@nelsonjedi can u explain how to do that? sorry...

OpenStudy (anonymous):

Let's do this one in steps: A) Subtract 9x from both sides, and you would have 0-3 = 0-3. Add 3 to both sides, and it is 0 = 6. B) Divide by 9, and you have x-3 = x-3. Again, 0 = 6. C) Subtract 9x from both sides, and you have -3 = 16x -3. 16x = -6, so x = -6/16. This one has a solution. D) Divide by 9, and x+3 = x+11. x = x+8, 0 = 8... Are you supposed to choose only one answer?

OpenStudy (anonymous):

yes isnt there two that i could choose?

OpenStudy (anonymous):

It appears that there should be 3 you can choose, A, B, and D, since a singular number cannot equal 6. Perhaps the question is, which one does have a solution? In which case, the answer would be C.

OpenStudy (anonymous):

I'm sorry, I meant a singular number cannot equal 0.

OpenStudy (anonymous):

@armyengineer225

OpenStudy (anonymous):

i dont know why it says that... there is a possibility it could of been a typo

OpenStudy (anonymous):

i went ahead and chose c it was wrong

OpenStudy (anonymous):

Perhaps, though I may be wrong. You may want to talk to your teacher.

OpenStudy (anonymous):

Ah I see, so they multiplied everything out.

OpenStudy (anonymous):

oh... okay

OpenStudy (anonymous):

Distribute to simplify and what do you have

OpenStudy (anonymous):

well can anybody help me with this? @armyengineer225 @luckyted @nelsonjedi

OpenStudy (anonymous):

solution to these equations: A) for every x in R, 9x - 3 = 9x -3, left side is exactly equals to right B) just like A. C) simplify to 9x=27x, so x = 0. D) no solution for x, such that 9x-27 = 9x-99. so the answer is D.

OpenStudy (anonymous):

oh that makes sense thank you so much @luckyted

OpenStudy (anonymous):

but i still need help with the last question i asked

OpenStudy (anonymous):

(x/5)+2 = 4. Subtract 2 from 4, you have (x/5) = 2. Multiply by 5 to clear the fraction, you have x = 10.

OpenStudy (anonymous):

thank you! @armyengineer225

OpenStudy (anonymous):

No problem. Sorry about the first question!

OpenStudy (anonymous):

it ok! :)